# The Nontrivial Vacuum Structure of an Extended $t\overline{t}$ BEH (Higgs) Bound state

### v3

@@ III.1 Formal Derivation
The action of eq.(10) is manifestly Lorentz invariant with dependence upon the 4-vector $\omega^{\mu}$. For the spherically symmetric ground state it must yield a Lorentz invariant expression in $\omega$. Hence, under a Lorentz transformation, $\omega^{\prime}_{\mu}=\Lambda_{\mu}^{\nu}\omega^{\nu}$ we can likewise perform $r^{\prime}_{\mu}=\Lambda_{\mu}^{\nu}r^{\nu}$, however, we are then free to change the integration variable back to its original form, $r^{\prime}\rightarrow r$ to obtain:
$\!\!\!\!\!\!\!\!\!\!S_{\phi}=M_{0}^{4}\int\!\!d^{4}r\biggl(Z\partial_{\mu}\phi_{\omega^{\prime}}^{\dagger}(r)\partial^{\mu}\phi_{\omega^{\prime}}(r)+2g_{0}^{2}N_{c}D_{F}(2r^{\mu})\phi_{\omega^{\prime}}^{\dagger}(r)\phi_{\omega^{\prime}}(r)\biggr).$ (21)
We can see that there is orthogonality of the $\phi_{\omega}(r^{\mu})$ solutions in the kinetic term (or in integrating over any extended smooth, approximately constant, Lorentz invariant function of $r^{\mu}$). The orthogonality breaks down, however, in the pointlike limit of the interaction where the $\phi_{\omega}(r)$ will freely mix with any $\phi_{\omega^{\prime}}(r)$ where $\omega^{\prime}\neq\omega$.
Formally, for a smooth Lorentz invariant function, $F(r^{\mu})$ (or differential operator, e.g, $F\sim\partial^{2}$):
@@
This implies that we will have the BCS or BEC-like behavior for $\Phi(r^{\mu})$. We derive the orthogonality in Appendix A.
Define a collective field by summing over a large set of $N$ arbitrary $\omega_{i\mu}$ unit 4-vectors that span the future timelike hyperboloid:
@@
The $\phi_{\omega_{i}}(r)$ each have an internal $\omega_{i\mu}$ and each is a solution of the equation of motion, eq.(11). We will use $\phi_{\omega}$ to represent any arbitrary $\omega$ field. In this framework the classical average of the $\phi_{\omega}$ is:
$\widetilde{\Phi}(r^{\mu})=\frac{1}{\sqrt{N}}\Phi(r^{\mu})=\frac{1}{{N}}\sum^{N}_{i=1}\phi_{\omega_{i}}(r^{\mu});\qquad\widetilde{\Phi}(r^{\mu})=\phi_{\omega_{i}}(0).$ (25)
Consider the action for the single two-body bound state $H(X^{\mu})\phi(r)$ of eqs.(81 to 85), replacing $\phi(r)\rightarrow\Phi(r^{\mu})$ and $H(x)\rightarrow H^{\prime}(X)$. and denote renormalized parameters by primes . For the sake of discussion we break the action into separate components:
$S=S_{1}+S_{2}+S_{3}+S_{Y}+S_{\lambda},\qquad\text{where,}\qquad H(X^{\mu})\phi(r)\rightarrow H^{\prime}(X)\Phi(r^{\mu}),$ (26)
where:
@@
and we have taken the pointlike limit of $D_{F}(2r^{\mu})$ of eq.(89),
@@
$\!\!\!\!\!\!\!\!\!\!S_{Y}=\widehat{g}^{\prime}_{Y}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{iL}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r)\;H^{\prime i}(X)\Phi(\vec{r}){+h.c.},\qquad\widehat{g}^{\prime}_{Y}\approx g^{\prime}{}_{0}^{2}\sqrt{2N_{c}/J},$ (31)
and the quartic interaction is given in the point-like limit of $\Phi(r^{\mu})$:
$\!\!\!\!\!\!\!\!\!\!S_{\lambda}=-\frac{\widehat{\lambda}}{2}\int\!\!d^{4}X(H^{\dagger}H)^{2}|\Phi(0)|^{4}{+h.c.}.$ (32)
Our problem is to verify that $S(H^{\prime}\Phi)$, with renormalized parameters, is consistent with the underlying theory of the $\phi_{\omega}=\phi(r)$ as defined in eqs.(81,87). We therefore substitute the collective field definition of eq.24 and define the renormalized parameters.
The orthogonality of the $\phi_{\omega_{i}}$ fields implies in a discrete sum:
We use these relations and the following normalizations:
@@
$H^{\prime}=H/\sqrt{N}\qquad\text{and then}\qquad Z^{\prime}={N}Z.$ (38)
@@
$=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|DH^{\prime}|^{2}\biggl(\frac{Z^{\prime}}{N}\sum^{N}_{i=1}\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{i}}(r)\biggr)=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|DH|^{2}\biggl(\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)\biggr)\;\text{for any }\omega,$
$=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|DH|^{2}|\widetilde{\Phi}|^{2}=\!\!\int\!\!d^{4}X|DH|^{2},$ (39)
where we have:
$ZM_{0}^{4}\int\!\!d^{4}r|\widetilde{\Phi}(r^{\mu})|^{2}=ZM_{0}^{4}\int\!\!d^{4}r|\phi_{\omega}(r^{\mu})|^{2}\;(\text{for any }\omega)=1.$ (40)
$S_{1}$ is therefore consistent with the underlying two-body theory. Note that $\widetilde{\Phi}_{\omega}$ and $\phi_{\omega}$ integrate out with the normalization eq.(40).
@@
$\!\!\!\!\!\!\!\!\!\!S_{2}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|H^{\prime}(X)|^{2}|\partial_{r}\Phi(r^{\mu})|^{2}\biggr)=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(|H|^{2}\partial_{\mu}\phi_{\omega}^{\dagger}(r)\partial^{\mu}\phi_{\omega}(r)\biggr)\;(\text{for any }\omega)$
$=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(|H|^{2}\partial_{\mu}\widetilde{\Phi}^{\dagger}(r)\partial^{\mu}\widetilde{\Phi}(r)\biggr),$ (41)
yielding consistency with the underlying kinetic term.
@@
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(2{g^{\prime}{}_{0}^{2}N_{c}}D_{F}(2r^{\mu})|H^{\prime}|^{2}|\Phi(r^{\mu}))|^{2}\biggr)\approx M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;{|H^{\prime}|^{2}}\biggl(2g^{\prime}{}_{0}^{2}N_{c}\frac{\delta^{4}(r^{\mu})}{16M_{0}^{2}}\frac{1}{N}\sum^{N}_{i=1}\phi_{\omega_{i}}^{\dagger}(r)\sum^{N}_{j=1}\phi_{\omega_{j}}(r)\biggr)$
$=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;N|H^{\prime}|^{2}\biggl(\frac{g^{\prime}{}_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)\phi^{\dagger}_{\omega}(0)\phi_{\omega}(0)\biggr)=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;|H|^{2}\biggl(\frac{g^{\prime}{}_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)\biggr)\;(\text{for any }\omega).$ (42)
The essential result is that the potential is approximately $\sim\delta^{4}(r^{\mu})$, hence the line integral orientation $\omega^{\mu}$ becomes irrelevant in the potential. In the second-to-last term we see the usual enhancement factor, $N$, that would normally lead to the BEC or BCS phenomena. However, in the last term we see that the normalization of $H^{\prime}=H/\sqrt{N}$ undoes the $N$-fold enhancement and the coupling is not renormalized $g_{0}^{\prime 2}=g_{0}^{2}$. This is therefore different than the case of the BCS superconductor or the BEC where the coupling is enhanced by a factor of $N$. The nonrenormalization of $g_{0}^{2}$ owes to the bilocal field theory with the renormalization of $H^{\prime}=H/\sqrt{N}$.
@@
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|H(X^{\mu})|^{2}\biggl(\frac{2g_{0}^{2}N_{c}}{M_{0}^{2}}D_{F}(2r)|\widetilde{\Phi}(r)|^{2}\biggr).$ (43)
To a good approximation we can freely swap between $D_{F}(2r)\leftrightarrow\delta^{4}(r)/16M_{0}^{2}$.
@@
$=g_{Y}M_{0}^{2}\int\!\!d^{4}X\;[\overline{\psi}_{iL}(X)\psi_{R}(X)]_{f}\;\;H^{i}(X)\widetilde{\Phi}(0){+h.c.};\qquad\qquad g_{Y}=g_{0}^{2}\sqrt{2N_{c}/J},$ (44)
where we have taken the limit $D_{F}(2r)\rightarrow\delta^{4}(r)/16M_{0}^{2}$.
@@
We see that $g_{Y}$ is nonrenormalized. The nonrenormalization of $g_{Y}$ implies that the result obtained for $M_{0}$ from the solution to the SKG equation remains intact, i.e., $M_{0}\sim 6$ TeV with $|\mu|\sim 88$ GeV. If we don’t take the strict $\delta$-function limit we can do an expansion of the integrand in $r^{\mu}$. This generates a series of higher dimension operators in inverse powers of $M_{0}^{2}$ providing potentially sensitive probes to $M_{0}$ and the shape of the wave-function $\Phi(r^{\mu})$, briefly described in section IV.
$S_{\lambda}$: The quartic term likewise involves,
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{1}{2}\hat{\lambda}^{\prime}\int d^{4}Xd^{4}r\;\delta^{4}(r)|H^{\prime}(X)\Phi(r^{\mu})|^{4}=\frac{1}{2}\hat{\lambda}\int d^{4}X|H(X^{\mu})|^{4}|\widetilde{\Phi}(0)|^{4},$ (45)
@@
where $\lambda^{\prime}=\lambda$ is therefore not renormalized. Essentially $\lambda$ is determined by the Yukawa coupling, $g_{Y}$ and the renormalization group running $\lambda$ from $M_{0}$ to $|\mu|$ and the result we obtained from the underlying theory at one loop, e.g., $\lambda\approx 0.23$, is not modified. This term would also permit expansion in $r^{\mu}$ and yield an operator expansion of new physics.
@@
Finally, the result of inserting solutions to the SKG equation, integrating the kinetic term by parts, yields the full action:
$S_{\Phi}=Z^{\prime}M_{0}^{4}\mu^{2}\int d^{4}Xd^{4}r|H^{\prime}|^{2}\;|\widetilde{\Phi}|^{2}=\mu^{2}\!\!\int d^{4}X|H|^{2}.$ (46)
@@
We thus obtain the same eigenvalue for $\Phi(r^{\mu})$ as for the underlying $\phi_{\omega}(r)$.
@@
It is straightforward to repeat the above with a continuous, manifestly Lorentz invariant, integral representation. To match definitions used above we have to reintroduce $N$, and require the matching condition to the classical sum:
$\sum_{i}^{N}(c)={\cal{N^{\prime}}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})\equiv\int_{\omega}(c)=N(c),$ (47)
where $c$ is an arbitrary constant. We then define the collective field, $\Phi(r^{\mu})$, and match the discrete and continuous representations,
$\Phi(r^{\mu})=\frac{1}{\sqrt{N}}\sum_{i}^{N}\phi_{\omega_{i}}={\cal{N}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})\equiv\frac{1}{\sqrt{N}}\int_{\omega}\phi_{\omega}(r^{\mu}).$ (48)
Hence ${\cal{N^{\prime}}}=\sqrt{N}{\cal{N}}$.
@@
$N={\cal{N}}\int\!\!d^{4}\hat{\omega}\;\delta(1-\hat{\omega}^{2})={\cal{N}}\pi^{2}\int\hat{\omega}^{2}d\hat{\omega}^{2}\delta(\hat{\omega}^{2}-1)=\pi^{2}{\cal{N}},$ (49)
@@
To apply this, consider the expression for the normalization of $H^{\prime}\Phi$, similar to the $S_{1}$ calculation above:
$Z^{\prime}M_{0}^{4}\int d^{4}Xd^{4}r|H^{\prime}|^{2}|\Phi(r^{\mu})|^{2}=Z^{\prime}M_{0}^{4}\int d^{4}Xd^{4}r|H^{\prime}|^{2}\frac{1}{\sqrt{N}}\int_{\omega}\phi^{\dagger}_{\omega}(r)\frac{1}{\sqrt{N}}\int_{\omega^{\prime}}\phi^{\dagger}_{\omega^{\prime}}(r)$
$=ZM_{0}^{4}\int d^{4}Xd^{4}r|H|^{2}\frac{1}{{N}}\int_{\omega}\phi^{\dagger}_{\omega}(r)\phi^{\dagger}_{\omega}(r)\times\biggl(\int_{\omega^{\prime}}\biggr)=ZM_{0}^{4}\int d^{4}Xd^{4}r|H|^{2},$ (50)
where we use the orthogonality relation,
$\int\!\!d^{4}r\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{j}}(r)=\delta_{ij}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)=\delta_{ij}\!\!\int d^{4}r|\widetilde{\Phi}(r^{\mu})|^{2},$ (51)
also eq.(47) for the dummy integral, $(\int_{\omega})=1$ and $Z^{\prime}|H^{\prime}|^{2}=Z|H|^{2}$ and eq.(40).
The interaction becomes:
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;|H^{\prime}|^{2}\biggl(2g^{\prime}{}_{0}^{2}N_{c}\frac{\delta^{4}(r^{\mu})}{16M_{0}^{2}}\frac{1}{\sqrt{N}}\int_{\omega}\phi_{\omega}^{\dagger}(r)\frac{1}{\sqrt{N}}\int_{\omega^{\prime}}\phi_{\omega^{\prime}}(r)\biggr)$
$=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;|H^{\prime}|^{2}\biggl(\frac{g^{\prime}{}_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)N\phi^{\dagger}_{\omega}(0)\phi_{\omega}(0)\biggr)=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;|H|^{2}\biggl(\frac{g_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)|\widetilde{\Phi}(r^{\mu})|^{2}\biggr).$ (52)

### v4

@@ III.1 Formal Derivation
The bound state action of natural top condensation in eq.(40) of ref.chris1 and takes the form:
where the composite field is $H(X^{\mu})\phi(r)$. The $H$ kinetic term is canonical with the Lorentz invariant normalization of eq.(13) and $Z$ is as defined upon removal of relative time as in eq.(12). The Yukawa interaction is generated at tree level,
In the above, the Yukawa and quartic couplings $g_{Y}$ and $\lambda$ are derived quantities from the underlying theory. In a pointlike approximation for the interaction, $D(2r)\sim\delta^{4}(r)/JM_{0}^{2}$, (where $J=16$ is the Jacobian for $(x,y)\rightarrow(X,r)$) and we have (from eq.(77) of ref.chris1 ):
$\widehat{g}_{Y}\approx g_{0}^{2}\sqrt{2N_{c}/J}\;\phi(0),\qquad\qquad\lambda\approx(g_{Y}^{4}-g_{Y}^{2}\lambda)\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{\mu}\right).$ (25)
Note that $\widehat{g}_{Y}$ is classical and $\lambda$ arises at loop level (${\cal{O}}(\hbar)$). We thus see that these are subject to dilution in an extended solution with the internal wave-function $\phi(r)$ and $g_{Y}=\widehat{g}_{Y}\phi(0)$ and $\phi(0)\propto\sqrt{|\mu|/M_{0}}$. Experimentally, with $|\mu|=88$ GeV, and inputting $g_{Y}\approx 1$ we determine $M_{0}\approx 6$ TeV, and $\lambda\approx 0.23$.
We now consider $\phi(r)\rightarrow\phi_{\omega}(r)$, the “Stueckelberg” fields of eq.(2) with relative time projected out. Our strategy is to replace, in the action, the composite field $H(X)\phi(r)$ considered in ref.chris1 , by the new field $H^{\prime}(X)\Phi(r)$, where $\Phi(r)$ is the coherent sum over $\phi_{\omega}(r)$. With suitable normaliziation conventions we will show that this reduces back to the standard model BEH action with the Lorentz invariant corrections of order $1/M_{0}^{2}$.
We can see that there is orthogonality between the $\phi_{\omega}(r^{\mu})$ solutions with different $\omega$’s. Formally, for a smooth Lorentz invariant function, $F(r^{\mu})$ (or differential operator, e.g, $F\sim\partial^{2}$):
@@
We derive the orthogonality in Appendix A.
For transparency, we will presently form a coherent state by a discrete sum over fields $\phi_{\omega_{i}}$.
$\!\!\!\!\!\!\!\!\!\!{\cal{N}}\int d^{4}\omega\delta(\omega^{2}-1)F(\phi_{\omega})\;\;\longrightarrow\;\;\sum^{N}_{i=1}F(\Phi_{\omega_{i}}).$ (28)
The $\phi_{\omega_{i}}(r)$ each have an internal $\omega_{i\mu}$ and each is a solution of the equation of motion, eq.(11). The orthogonality of the $\phi_{\omega_{i}}$ fields implies in a discrete sum:
Therefore, we define the collective field by summing over a large set of $N$ arbitrary $\omega_{i\mu}$ unit 4-vectors that span the future timelike hyperboloid:
@@
In this framework we will also have the classical average of the $\phi_{\omega}$:
$\widetilde{\Phi}(r^{\mu})=\frac{1}{\sqrt{N}}\Phi(r^{\mu})=\frac{1}{{N}}\sum^{N}_{i=1}\phi_{\omega_{i}}(r^{\mu});\qquad\text{Note that for any $\omega_{i}$:}\qquad\widetilde{\Phi}(0)=\phi_{\omega_{i}}(0).$ (31)
Consider the action for the single two-body bound state $H(X^{\mu})\phi(r)$, replacing $\phi(r)\rightarrow\Phi(r^{\mu})$ and $H(x)\rightarrow H^{\prime}(X)$. and denote renormalized parameters by primes . For the sake of discussion we break the action into separate components:
$S=S_{1}+S_{2}+S_{3}+S_{Y}+S_{\lambda},\qquad\text{where we define,}\qquad H(X^{\mu})\phi(r)\rightarrow H^{\prime}(X)\Phi(r^{\mu}),$ (32)
and:
@@
where we have here taken the pointlike limit of $D_{F}(2r^{\mu})$ of eq.(90),
@@
$\!\!\!\!\!\!\!\!\!\!S_{Y}=\widehat{g}_{Y}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{iL}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r)\;H^{\prime i}(X)\Phi(\vec{r}){+h.c.},$ (37)
and the quartic interaction,:
$\!\!\!\!\!\!\!\!\!\!S_{\lambda}=-\frac{\widehat{\lambda}}{2}\int\!\!d^{4}X(H^{\prime\dagger}H^{\prime})^{2}|\Phi(0)|^{4}{+h.c.}.$ (38)
Our problem is to verify that $S(H^{\prime}\Phi)$, with renormalized parameters, is consistent with the underlying theory of the $\phi_{\omega}=\phi(r)$ as defined in eqs.(32,88). We therefore substitute the collective field definition of eq.30 and define the renormalized parameters.
We define the following normalizations:
@@
$H^{\prime}=H/\sqrt{N},\qquad\text{and then,}\qquad Z^{\prime}={N}Z.$ (43)
@@
$=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|DH^{\prime}|^{2}\biggl(\frac{Z^{\prime}}{N}\sum^{N}_{i=1}\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{i}}(r)\biggr)=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|DH|^{2}\biggl(\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)\biggr)\;\text{(for any }\omega),$
$=\!\!\int\!\!d^{4}X|DH|^{2},$ (44)
where we use the normalization of the $\phi_{\omega_{i}}$ as in eq.(15).
@@
$\!\!\!\!\!\!\!\!\!\!S_{2}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|H^{\prime}(X)|^{2}|\partial_{r}\Phi(r^{\mu})|^{2}\biggr)=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(|H|^{2}\partial_{\mu}\phi_{\omega}^{\dagger}(r)\partial^{\mu}\phi_{\omega}(r)\biggr)\;\;\;(\text{for any }\omega),$ (45)
yielding consistency with the underlying kinetic term of the action eq.(22).
@@
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(2{g{}_{0}^{2}N_{c}}D_{F}(2r^{\mu})|H^{\prime}|^{2}|\Phi(r^{\mu}))|^{2}\biggr)\approx M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;{|H^{\prime}|^{2}}\biggl(2g{}_{0}^{2}N_{c}\frac{\delta^{4}(r^{\mu})}{16M_{0}^{2}}\frac{1}{N}\sum^{N}_{i=1}\phi_{\omega_{i}}^{\dagger}(r)\sum^{N}_{j=1}\phi_{\omega_{j}}(r)\biggr)$
$=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;N|H^{\prime}|^{2}\biggl(\frac{g{}_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)\phi^{\dagger}_{\omega}(0)\phi_{\omega}(0)\biggr)=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;|H|^{2}\biggl(\frac{g{}_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)\biggr)\;(\text{for any }\omega).$ (46)
The essential result is that the potential is approximately $\sim\delta^{4}(r^{\mu})$, hence the line integral orientation $\omega^{\mu}$ becomes irrelevant in the potential. In the second-to-last term we see the usual enhancement factor, $N$, that would normally lead to the BCS enhancement phenomena. However, in the last term the normalization of $H^{\prime}=H/\sqrt{N}$ undoes the $N$-fold enhancement and the coupling $g_{0}^{2}$ is not renormalized. The nonrenormalization of $g_{0}^{2}$ owes to the bilocal field theory with the renormalization of $H^{\prime}=H/\sqrt{N}$.
@@
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|H(X^{\mu})|^{2}\biggl({2g_{0}^{2}N_{c}}D_{F}(2r)|\widetilde{\Phi}(r)|^{2}\biggr).$ (47)
In the approximation $\sim{\cal{O}}(1/M_{0}^{2})$ we can freely swap between $D_{F}(2r)\leftrightarrow\delta^{4}(r)/16M_{0}^{2}$, and here we use eq.(30) where $\widetilde{\Phi}(0)=\phi_{\omega}(0)$ for any $\omega$.
@@
$=\widehat{g}_{Y}M_{0}^{2}\int\!\!d^{4}X\;[\overline{\psi}_{iL}(X)\psi_{R}(X)]_{f}\;\;H^{i}(X)\widetilde{\Phi}(0){+h.c.};\qquad\qquad(\widehat{g}_{Y}=g_{0}^{2}\sqrt{2N_{c}/J}),$ (48)
where we have taken the limit $D_{F}(2r)\rightarrow\delta^{4}(r)/16M_{0}^{2}$, and the field $\widetilde{\Phi}(r)$ appears here. Then $\widehat{g}_{Y}\widetilde{\Phi}(0)=\widehat{g}_{Y}\phi_{\omega}(0)=g_{Y}$, hence we revert to the standard model:
$\!\!\!\!\!\!\!\!\!\!S_{Y}\rightarrow{g}_{Y}M_{0}^{2}\int\!\!d^{4}X\;[\overline{\psi}_{iL}(X)\psi_{R}(X)]_{f}\;H^{i}(X){+h.c.}.$ (49)
@@
We see that $g_{Y}$ is nonrenormalized. Since the classical average field, $\widetilde{\Phi}(0)=\phi_{\omega}(0)$, appears here it imples that the result obtained for $M_{0}$ from the solution to the SKG equation remains intact, i.e., $M_{0}\sim 6$ TeV with $|\mu|\sim 88$ GeV. If we don’t take the strict $\delta$-function limit we can do an expansion of the integrand in $r^{\mu}$. This generates a series of higher dimension operators in inverse powers of $M_{0}^{2}$ providing potentially sensitive probes to $M_{0}$ and the shape of the wave-function $\Phi(r^{\mu})$, briefly described below in section IV.
$S_{\lambda}$: The quartic term likewise becomes,
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{\lambda}=-\frac{1}{2}\hat{\lambda}\int d^{4}Xd^{4}r\;\delta^{4}(r)|H^{\prime}(X)\Phi(r^{\mu})|^{4}=-\frac{1}{2}\hat{\lambda}\int d^{4}X|H(X^{\mu})|^{4}|\widetilde{\Phi}(0)|^{4}=-\frac{1}{2}{\lambda}\int d^{4}X|H(X^{\mu})|^{4},$ (50)
@@
where $\lambda=\hat{\lambda}|\widetilde{\Phi}(0)|^{4}$ is therefore not renormalized. Essentially $\lambda$ is determined by the Yukawa coupling, $g^{4}_{Y}$, through the usual renormalization group running $\lambda$ from $M_{0}$ to $|\mu|$ and the successful result obtained from the underlying theory at one loop, e.g., $\lambda\approx 0.23$, is not modified. This term would also permit an expansion in $r^{\mu}$ and yield an operator expansion of new physics.
@@
Finally, the result of inserting solutions to the SKG equation, integrating the kinetic term by parts, yields the eigenvalue of the full action:
$S_{\Phi}=-Z^{\prime}M_{0}^{4}\mu^{2}\int d^{4}Xd^{4}r|H^{\prime}|^{2}\;|{\Phi}|^{2}=-\mu^{2}\!\!\int d^{4}X|H|^{2},$ (51)
where we have with $Z^{\prime}H^{\prime}=ZH$, and, for any $\omega_{i}$:
$ZM_{0}^{4}\int\!\!d^{4}r|\Phi(r^{\mu})|^{2}=ZM_{0}^{4}\int\!\!d^{4}r|\phi_{\omega_{i}}(r^{\mu})|^{2}=1,$ (52)
@@
We thus obtain the same eigenvalue for the action with $\Phi(r^{\mu})$ as for the action with the underlying $\phi_{\omega}(r)$.
@@
It is straightforward to repeat the above with a continuous, manifestly Lorentz invariant, integral representation. To match definitions used above we require the matching condition to the classical collective field, $\Phi(r^{\mu})$, betweeb the discrete and continuous representations,
$\Phi(r^{\mu})=\frac{1}{\sqrt{N}}\sum_{i}^{N}\phi_{\omega_{i}}={\cal{N}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})\equiv{\cal{N}}\int_{\omega}\;\phi_{\omega}(r^{\mu}).$ (53)
To apply this, consider the expression for the normalization of $\Phi$ above:
$1=ZM_{0}^{4}\int d^{4}r|\Phi(r^{\mu})|^{2}=ZM_{0}^{4}\int d^{4}r\;{\cal{N}}\int_{\omega}\phi^{\dagger}_{\omega}(r)\;{\cal{N}}\int_{\omega^{\prime}}\phi_{\omega^{\prime}}(r)$
$=ZM_{0}^{4}\int d^{4}r\;{\cal{N}}^{2}\int_{\omega}\phi^{\dagger}_{\omega}(r)\phi^{\dagger}_{\omega}(r)=ZM_{0}^{4}\int d^{4}r\frac{1}{N}\sum\phi^{\dagger}_{\omega}(r)\phi_{\omega}(r),$ (54)
where we use the orthogonality relation, eq.(26), hence we have: $1={\cal{N}}^{2}N$.
@@
${\cal{N}}\int\!\!d^{4}\hat{\omega}\;\delta(1-\hat{\omega}^{2})={\cal{N}}\pi^{2}\int\hat{\omega}^{2}d\hat{\omega}^{2}\delta(\hat{\omega}^{2}-1)=\pi^{2}{\cal{N}},$ (55)